If z varies inversely as​ w, and z=30 when w=0.4​, find z when w=10.
Answer by Alan3354(69443)

If z varies inversely as​ w, and z=30 when w=0.4​, find z when w=10.
------------
z = k/w --- k is a constant
30 = k/0.4
k = 12
==================
find z when w=10.
z = 12/10
z = 1.2